A flat metal belt made of Beryllium-Copper (BeCu) is Sf = 375 MPa and transmits at 8.24 kW / min and 8.25 kW of power. The belt life is 935 mm and 106 turns. The friction coefficient between the belt and the pulley will be 0.34, the operating coefficient is 1.41, and the design coefficient will be 1.55. Under the rotated pulley, the motor is 3.25 times the pulley diameter. Calculate belt belt and length for 0.20 mm in belt area and for pulley with minimum diameter. Belt width is required in 5 mm and its layers. Calculate the pretension force to be applied.
Answer #1

solution vriven, Sp=345MPa Becu Transmitted Power PL) = 8.25 KW co-efficient of friction (A) =0-34 Operating co-efficient (kgLength of y flat belf (L) = [ue==(2-doba \(8@$+8,8%) -- [uryeso?- (16 29-10612_) 0-4 123 vastasy 805) L=2766-75 mm Length of